# Interpreting Shared Circuits for Ordered Sequence Prediction in a Large Language Model

### v3

@@ 5.2.2 Last Token Sequence Detection Head
Figure 4: Attention Pattern for Head 7.11. At the last token, head 7.11 to more recent sequence members than earlier sequence members. We also recognize an offset pattern on the diagonals of this heatmap, indicating that it also functions as a previous token head. Note we take the mean of from the Numerals dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration.
Figure 5: The Attention Pattern of Head 9.1 shows the component pays strong attention to only the most recent number. Note we take the mean of from the Numerals dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration.
5.2.3 Next Sequence Head
Figure 2 shows that head 9.1 receives information from both head 4.4 and 7.11. Head 9.1, shown in Figure 5, pays strong attention to only the last member of the sequence, and it appears to attend even stronger to the last member than 7.11.
Next Sequence Scores. To check that head 9.1 outputs next sequence tokens, we study its component output scores. Table 3 shows that given a numeral token I as input (eg. 1), head 9.1 often outputs a token I+1 or higher (eg. 2). For numerals between 1 and 100, its next score is 87%, while its copy score is 59%. We also note that the next sequence scores of most heads are low, with an average of 3.29%, and that head 9.1 has the highest next sequence score. Thus, it seems to function as a "next sequence head".  [^8] In a case of multiple independent discovery, successor heads were also found by Gould et al. (2023). This is reinforced by the next sequence score for number words, which is 90.63%, while the average for all attention heads is 2.97%. Although head 9.1 does not appear to output months given any month token, we observe something peculiar: 9.1 is the only head that will output the next rank given a month (eg. given "February", output "third", and its "next rank given month" score is 31.25%. This appears to be related to how months can be mapped onto ranks.
@@ 5.2.3 Next Sequence Head
5.3 MLP Functionality
For most samples of all sequence types, logit lens reveals that the model does not predict the correct answer before MLP 9. However, after the information is processed through MLP 9, the model outputs the next sequence member. These findings suggests that MLP 9 is largely responsible for finding the next sequence member, even more so than Head 9.1, which may just be boosting this information and/or acting as a backup component. Logit lens results can be found in Appendix §E.
6 Conclusion
Understanding the inner workings of neural networks is essential for fostering alignment and safety. In this paper, we identify that across similar sequence continuation tasks, there exist shared sub-circuits that exhibit similar functionality. Specifically, we find sequence token detection heads and components associated with next sequence outputs. The aim of this work is to advance our understanding of how transformers leverage shared computational structures across similar tasks. By locating and comparing these circuits, we hope to gain insight into semantic representations of abstract concepts, which may provide evidence for hierarchical associations.
7 Limitations
Number of Models. As it is common for many well-received interpretability papers to focus on analyzing one model Wang et al. (2022); Conmy et al. (2023); Hanna et al. (2023); Nanda et al. (2023) or even just one attention head (McDougall et al., 2023), we only analyze one model. We plan to investigate this phenomenon for multiple models, including toy models and larger models, that can perform other types of sequence continuation in future work. This future work may include Fibonacci sequences, or comparing circuits for adding 2 (e.g. 2 4 6 8 10) vs circuits for multiplying 2 (e.g. 2 4 8 16 24), studying if the model uses diverging circuits to differentiate between the two tasks while still computing parts of them using shared circuits.
Dataset Size. As there are only twelve months, the number of possible continuing sequences was limited. Additionally, even if months were not used, GPT-2 also has limited prediction ability for number word sequences, as detailed in Appendix B. However, our dataset size for the months continuation task fully captures all of the months; in contrast, a small dataset size brings more issues when it does not capture all of the true distribution.
Task Complexity. Though sequence continuation may be seen as possibly simply associating what comes after every sequence member, the aim of our work is to find how this task is internally represented, even if it is done by simple association. Previous works have investigated how association is done by MLPs (Geva et al., 2020), so understanding how language models associate information can shed light on how it represents similar concepts.
Appendix A All Three Circuits Diagram
Appendix B Dataset Details
@@
B.1 IOI Circuit.
The IOI circuit we use for comparison in Table 1 uses all MLPs and the following heads:
(0, 1), (0, 10), (2, 2), (3, 0), (4, 11), (5, 5), (5, 8), (5, 9), (6, 9),
(7, 3), (7, 9), (8, 6), (8, 10), (9, 0), (9, 6), (9, 7), (9, 9),
(10, 0), (10, 1), (10, 2), (10, 6), (10, 7), (10, 10), (11, 2), (11, 9), (11, 10)
Appendix C Computational Resources and Packages
Appendix D Individual Circuit Results
@@
Appendix E MLP Analysis Details
@@ 11 September, August, October
Appendix F Iterative Method Details
Instead of absolute impact on performance score, we can also use relative impact. For instance, this means that the removal won’t cause the score to go down by more than 0.01 of the existing score, rather than an absolute threshold of 80%. However, this still doesn’t take combinations into account; one edge removal may decrease the score by 0.01, and another may decrease it by 0.01, but when both are removed, their combined effect may not make the model score go down by 0.02; it may go down by more. Thus, the order of component and edge removal appears to an impact on the circuit that is found.
Appendix G Functionality Method Details
@@
Appendix H Attention Pattern Extended Results
@@
Appendix I Months Circuit Keeping MLP 11
Although the iterative node pruning algorithm removes MLP 11 for the Months task, we note this is only because the performance drop of -19.50%, shown in Table 6, is barely within threshold $T=20\%$. Thus, we also run experiments to iteratively find a Months circuit that keeps MLP 11, which is shown with the other two tasks’ circuits in Figure 13. Table 11 shows the importance of each head in the circuit. Overall, the results are largely similar to the Months circuit shown in Figure 2.
Figure 13: Showing all three circuits and their overlap, but using a Months circuit that keeps MLP 11.
Table 11: Performance Drop when Head is Removed from Months Circuit with MLP 11.
Important Head Months
(4, 4) -78.27
(7, 11) -68.61
(0, 1) -61.09
(9, 1) -51.60
(6, 10) -33.00
(8, 8) -12.47
(6, 6) -11.87
(4, 10) -6.95
(7, 10) -4.74
(6, 1) -4.52
(7, 2) -2.06
Appendix J Circuit Entanglement Definitions
Definitions. A circuit can be defined as "a human-comprehensible subgraph, which is dedicated to performing certain task(s), of a neural network model" (Räuker et al., 2023). To describe the circuits representations in this paper, we define a circuit graph as a connected graph $C$ with (1) a node set $N$ of components, and (2) an edge set $E$, in which an edge $(n_{1},n_{2})$ represents how component $n_{1}$ affects of component $n_{2}$.  [^9] Different studies may define ”circuit” in different ways. We define that a circuit graph $C$ is used for a task $T$ based on how ablating all model components aside from those in the circuit still allows the model to have a certain level of performance; we determine this level as described in iterative pruning in §3.1. We note that a task $T$ can be broken into a set of sub-tasks $S_{T}=\{S_{1},...,S_{n}\}$; for instance, one sub-task of IOI is to inhibit repeated subjects. Next, we define a circuit graph $C_{1}$ to be a circuit subset of circuit graph $C_{2}$ if all the nodes in the node set of $C_{1}$ are contained in the set of nodes for $C_{2}$. We also define $C_{1}$ to be a sub-circuit of $C_{2}$ all the edges in the edge set of $C_{1}$ are also contained in the edge set of $C_{2}$. We further define circuit graph $C_{1}$ used for task $T_{1}$ to be a functional sub-circuit (or subset) of circuit graph $C_{2}$ used for task $T_{2}$ if these conditions are met: (1) $C_{1}$ is a sub-circuit of $C_{2}$, and (2) $T_{1}$ is a subtask in $S_{T_{2}}$, the set of subtasks of $T_{2}$.
Circuit Entanglement. Given that components play multiple roles (Merullo et al., 2023), editing components in a circuit used for task A can have an effect on task B. Instead of just vaguely assuming this would have "some effect", such as ruining task B in "some way", our aim is to precisely describe, and thus approximately predict, what this effect is. We define two circuits $C_{1}$ and $C_{2}$ as being analogously entangled if editing the functionality of $C_{1}$ affects the functionality of $C_{2}$ in an analogous way. For instance, let $C_{1}$ be for "digits continuation" and let $C_{2}$ be for "months continuation". If component $H$ in $C_{1}$ finds the "next digit of a sequence" and we edit it to now find the "previous digit", then if task B now finds the "previous month", we say $C_{1}$ and $C_{2}$ are analogously entangled.
Appendix K Broader Impact Statement

### v4

@@ C.1 Last Token Sequence Detection Head
C.2 Successor Head Scores
To check that head 9.1 outputs next sequence tokens, we study its component output scores. Table 5 shows that given a numeral token I as input (eg. 1), head 9.1 often outputs a token I+1 or higher (eg. 2). For numerals between 1 and 100, its next score is 87%, while its copy score is 59%. We also note that the next sequence scores of most heads are low, with an average of 3.29%, and that head 9.1 has the highest next sequence score. Thus, it seems to function as a "next sequence head".
This is reinforced by the next sequence score for number words, which is 90.63%, while the average for all attention heads is 2.97%. Although head 9.1 does not appear to output months given any month token, we observe something peculiar: 9.1 is the only head that will output the next rank given a month (eg. given "February", output "third", and its "next rank given month" score is 31.25%. This appears to be related to how months can be mapped onto ranks.
@@ C.2 Successor Head Scores
Appendix D Llama-2-7B Functionality Details
As the full circuits are too large to include in this paper, we only discuss important components.
Sequence Continuation: Other Attention Patterns Figures 7 and 8 show attention patterns for heads 11.8 and 16.0, which were found to be common in all three sequence continuation tasks. These have a slight resemblence to the "last token sequence detection head" 7.11 in GPT-2 Small, though the resemblence is not as prominent as the mapping between 4.4 and 5.25, and 9.1 and 20.17, in GPT-2 Small vs Llama-2-7B. Despite head 16.0 being more important than head 11.8, its attention pattern is not as strong.
Figure 6: Llama-2-7B attention pattern for Head 20.17, bearing resemblence to the attention pattern of GPT-2 Small’s Head 9.1 in Figure 3, showing the component pays strong attention to only the most recent sequence member.
Figure 7: Llama-2-7B attention pattern for Head 11.8. At the last token, there appears to be a resemblence to head 7.11, where more recent sequence members than earlier sequence members.
Figure 8: Llama-2-7B attention pattern for Head 16.0. Tokens appears to attend to the last period of the first instance of the pattern (<object> lost in Month.) However, it is not clear from this plot what the role of this head is. Despite being more important than head 11.8, its attention pattern is not as strong.
Interval-K Sequences: Attention Patterns We apply the iterative ablation methods to intervaled sequences such as additive +2 (eg. "2 4 6"). We discover that heads 16.0 and 20.17 are within these components sets, and that head 6.11 is considered especially important. Future studies can delve more into this area.
The attention patterns for the important attention heads 5.25 and 20.17 on the Interval-K tasks, where $K\geq 1$, are similar to those for Interval-1. For intervals that are not a multiple of 10 and for numbers with more than one digit, the last digit of a member attends to the last digit of a previous member. For intervals that are a multiple of 10, these digits attend to the differing number (eg. the ’3’ in 300 attends to the ’2’ in 200).
Appendix E Computational Resources and Packages
Appendix F GPT-2 Experimental Setup Details
Circuits Variations. We observe that there are multiple circuits, with slight variations between them, that have similar performances for the same task. However, we find that important heads are often found in most circuits, regardless of the method, metric or dataset choices. Thus, we focus more on the "big picture" comparison of scores and on the most important heads, and less on the exact variations between scores or the less important heads.
@@ Appendix F GPT-2 Experimental Setup Details
F.1 IOI Circuit.
The IOI circuit in Table 4 uses all MLPs and the following heads (shown in (L,H) format):
(0, 1), (0, 10), (2, 2), (3, 0), (4, 11), (5, 5), (5, 8), (5, 9), (6, 9), (7, 3), (7, 9), (8, 6), (8, 10), (9, 0), (9, 6), (9, 7), (9, 9), (10, 0), (10, 1), (10, 2), (10, 6), (10, 7), (10, 10), (11, 2), (11, 9), (11, 10)
Appendix G Llama-2 Experimental Setup Details
Dataset. For Llama-2, we ran iterative node pruning on samples consisting only of sequence members, and ran attention patterns on samples consisting of both sequence and non-sequence members. This is because for Llama-2, the main focus was on finding similar components between the sequence continuation sub-circuits of GPT-2 and Llama-2, rather than finding more minimal circuits, which would be beyond the computational resources we allocated.
Iterative Methods. Given that Llama-2 is larger than GPT-2, we modified the iterative component ablation method to first ablate all the heads of an attention layer to see if this makes the model performance perform below the threshold; if it does, we remove all the heads of that attention layer from the candidate circuit. While this method variation is faster if many attention layers are unnecessary, it has the downside in which removing certain unimportant heads may increase performance, which may, in summation, offset the performance degrading abilities of important heads. This means that there may be some important heads which may be missed by this method. However, we found that the most important attention heads were not missed by this method, and the end results were often comparable to the original version of our iterative methods.
Appendix H All Three Circuits Diagram for GPT-2 Small
Appendix I Individual Circuit Results for GPT-2 Small
@@
Appendix J MLP Analysis Details for GPT-2 Small
@@ 11 September, August, October
Appendix K Functionality Method Details
@@
Appendix L Attention Pattern Extended Results for GPT-2 Small
@@
Appendix M Broader Impact Statement